OpenStudy (anonymous):

Find the equation of a circle that has a diameter with endpoints at (0, 4) and (-3, 1) then tell me how you did it :)

OpenStudy (anonymous):

circle with a diameter? no way

OpenStudy (amistre64):

lol .... determine mid point

OpenStudy (anonymous):

What's the midpoint formual again :D?

OpenStudy (amistre64):

hlafway

OpenStudy (amistre64):

halfway

OpenStudy (amistre64):

add the points and divide by 2

OpenStudy (anonymous):

uh.. okay

OpenStudy (anonymous):

(0,4) + (-3,1)?

OpenStudy (amistre64):

(0, 4) (-3, 1) ------ (-3 , 5) /2 = (-3/2, 5/2)

OpenStudy (amistre64):

for example; whats halfway between 6 and 8? 6+8 = 14, /2 = 7 right?

OpenStudy (anonymous):

sure now that I have that, what can I do with it?

OpenStudy (amistre64):

use it for the center :)

OpenStudy (amistre64):

now determine the distance between your points by subtracting them and pythaging it

OpenStudy (amistre64):

(0, 4) -(-3, 1) ------- 3 , 3 ; well, thats good; its just 3 sqrt(2)

OpenStudy (anonymous):

pythaging it?

OpenStudy (anonymous):

isn't their an easier, more algebraic way of finding the equation here?

OpenStudy (amistre64):

yeah; sqrt(x^2 + y^2) ... the pythag thrm

OpenStudy (anonymous):

there* wow

OpenStudy (amistre64):

yeah, get someone else to do it :)

OpenStudy (amistre64):

but no, this is the nuts and bolts of it

OpenStudy (amistre64):

since we now the distance from point to point; split it in half to determine the radius

OpenStudy (amistre64):

and put it all together now: (x+(3/2))^2 + (y-(5/2))^2 = (3sqrt(2)/2)^2